OpenStudy (anonymous):

How do I complete fractions? 3/7 + 1/7

OpenStudy (anonymous):

(3+1)/7=4/7

OpenStudy (anonymous):

same the denumerator, and suit with numeratot

OpenStudy (anonymous):

Thanks infauzan!

OpenStudy (anonymous):

\[\color{blue}{\text{denumerator!}}\]

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

And.... 1/6 x 3/4 And.... 5/8 - 1/2 And.... 6/7 divided by 3/14

OpenStudy (anonymous):

\[\frac{1}{6}\times \frac{3}{4}=\frac{1\times 3}{6\times 4}=\frac{1}{2\times 4}=\frac{1}{8}\]

OpenStudy (anonymous):

1. 1/4x3/6=1/8 2. (5-4)/8=1/8 3.6 /7x14/3=6/3x14/7=4

OpenStudy (anonymous):

make it simple

OpenStudy (anonymous):

\[\frac{6}{7}\div\frac{3}{14}=\frac{6}{7}\times \frac{14}{3}=2\times 2=4\]

OpenStudy (anonymous):

Thanks satellite73 and infauzan!